OpenStudy (anonymous):

show that the area of a circle x^2+y^2=a^2 is pi a^2

OpenStudy (shubhamsrg):

what is the radius of the circle : x^2 + y^2 = a^2 ??

OpenStudy (anonymous):

solve for y and get y = sqrt(a^2-x^2) <- equation for a semi circle integrate this from -a to a  once you integrate it you will get \[\Pi a^2 / 2 \] since you found the area of a semicircle multiply it by 2 to get the area of the circle which is \[\Pi a^2\]